# Tokens, the oft-overlooked appetizer: Large language models, the distributional hypothesis, and meaning

### v8

@@ 4 Conclusion
It is clear that we need to develop strategies that allow LLMs to make better use of the data to which they are exposed.109109109Zhuang et al., 2024 note “the immense volume of training data necessitated for effective LM performance, surpassing—by orders of magnitude—the linguistic input received during human language acquisition” Zhuang, Fedorenko, and Andreas (2024). People, by contrast, are very good at wringing new information from familiar stories.110110110Right now, models display a striking dissimilarity with people in what happens when they are repeatedly exposed to the same data, either directly, or indirectly through iterated training on synthetic data. For example, Chang et al., 2024 found that “LLMs trained with duplicated training data exhibit faster forgetting” Chang et al. (2024), and Shumailov et al., 2024, found “indiscriminate use of model-generated content in training causes irreversible defects in the resulting models, in which tails of the original content distribution disappear” Shumailov et al. (2024). On the other hand, a hallmark of human storytelling is enjoying—and getting new content out of—experiencing a story (created by other humans, even potentially by oneself) again and again. People have been studying religious texts, as in Biblical exegesis, for thousands of years, yet new ideas emerge. An individual may return to the Bible (continuing the same example) upwards of hundreds of times, and yet experience new insight. How can LLMs wring new insight from old data? This stark difference in behaviour is worth further consideration.
@@
One strategy would be to bring various representations the LLM uses into greater alignment with human structures and processes. There is evidence that humans can do quite a lot with very little data (e.g., a broad spectrum of poverty-of-the-stimulus-related arguments and processes Zhuang, Fedorenko, and Andreas (2024), including fast-mapping Gelman and Brandone (2010), and in experimental scenarios testing the distributional hypothesis McDonald and Ramscar (2001)).111111111Likewise, LLMs can sometimes do a lot with a little, such as in lexinvariant LMs Huang et al. (2023b).
@@
Furthermore, good model vocabularies may already be converging on being somewhat human-like. Symmetrically, the tokens themselves point to commonalities between the entropy-minimizing objective functions often used in tokenization and the forces that shape human language. The hundreds of thousands of years Fedorenko, Piantadosi, and Gibson (2024)112112112E. Fedorenko is an author on both Fedorenko et al.,, 2024 Fedorenko, Piantadosi, and Gibson (2024) and Zhuang et al.,, 2024 Zhuang, Fedorenko, and Andreas (2024), which seem to make arguments that are at least partially contradictory. We assume in good faith there must be a perspective (presumably E.F.’s) from which the arguments are entirely consistent, and take both papers at face value for the purposes of this project. in which only people spoke language, during which pressures of efficiency seem to have shaped its development at multiple levels, surely evolved structures we can now take advantage of in our models.
@@
We argue that the de facto approaches in each of those areas are worth revisiting.113113113The use of special tokens (e.g., SEP) in LLMs is out of scope for this paper but intriguing for future work, as is the impact of embedding vector dimension and other implementation choices. More broadly, now that the efficacy of LLM architecture has been established, we should revisit our prior underconstrained implementation choices, such as the relevant objective functions that shape the model and the functions that conjoin its components.114114114e.g., operations involving Query, Key, and Value vectors in the attention blocks, or applications of softmax.
@@
We contend that LLMs validate and isolate the distributional hypothesis. Specifically, they show the power of syntagmatic information for producing minimally viable language performance (when coupled with sufficient architecture). We hypothesize that LLMs bootstrap limited meaning, potentially including abstractions and world knowledge, from distributional patterns: more specifically, that while LLM’s start with syntagmatics, they create further structure leading to limited semantic meaning, the ceiling of which is what can be conveyed by purely distributional linguistic information and diegetic approximations composed thereof.115115115Harris’ purely linguistic meaning Sahlgren (2008). The lack of hard boundary between world knowledge and linguistic knowledge makes this feat possible but it also engenders the problems that come from fuzzing up grammaticality and truth, or frequency and truth Zimmerman et al. (2024); Gong et al. (2020),116116116Gong et al., 2020, found that in some situations, word embeddings were “biased towards word frequency”, with the embeddings of high- and low-frequency words lying in different areas of the embedding space, and the embedding of a rare word and a popular word being physically distant despite their semantic similarity. Within the embedding space, rare and popular words behaved differently, with popular words having semantically relevant neighbors, and rare words being overwhelmingly surrounded by other, semantically-unrelated, rare words Gong et al. (2020). (They foresee these structures as causing downstream performance issues, so they propose FRequency AGnostic word Embedding (FRAGE) as a solution Gong et al. (2020).) which are amplified by the choice of next token prediction as objective function.117117117And by other, more specific, architectural choices, such as how attention is implemented. These choices are related, since the attention design is in the service of generating plausible output tokens, so in the context of this paper, we are referring to the broad goal of getting the model to generate plausible tokens, not only to specific portions of the model that might be most directly technically implicated, like the decoder (if there is one), the output layer, or the final probability distribution over the vocabulary. We propose that experiential grounding enhances the depth of meaning through othermodal reference for both people and LLMs118118118Recall Wolfe and Caliskan, 2022 Wolfe and Caliskan (2022) and Zhuang et al., 2024 Zhuang, Fedorenko, and Andreas (2024). and therefore represents a promising avenue for their continued improvement.
@@
Importantly, a model’s gnogeography119119119Recall that gnogeography is the landscape of knowledge; a being’s mind-and-body, their architecture, in conjunction with what information they have taken in. can be expanded through diegetic approximations Zimmerman et al. (2024); Brown et al. (2020); Chang et al. (2024); Mahowald et al. (2024). Meaning that, with careful structuring of training data and model architecture (e.g., batch size, number of training steps), LLMs do seem to be able to acquire additional world knowledge Chang et al. (2024), implying that there is space between the world knowledge necessary for MVP language and what it is possible to know through the medium of language. Given that current LLMs clearly possess MVP language but are missing much human-familiar knowledge and common sense, they are somewhere within that space Zimmerman et al. (2024); Mahowald et al. (2024). The entirety of that space is the upper bound of their gnogeography.
This is crucial to remember in the face of AI hype, one core promise of which is “we have a machine that can talk, therefore we have a machine that can think”. In fact, LLMs demonstrate the plausibility of a machine that can solely talk; while a machine may be able to both talk and think, the ‘therefore’ between those abilities is not justified.120120120“Talk” and “think” are shorthand. Of course, talking involves thinking, and thinking can be scaffolded by talking. See previous sections, especially Sec. 2.3, where we referenced stochastic parrots; we do think MVP language as implemented in LLMs involves more understanding than would be granted by that framework.
The size and shape of the upper bound LLM gnogeography is unknown.121121121The hypothetical ceiling being everything that can be encoded in language. While this is exciting, it also means it is plausible that their gnogeography is incompatible with what would be required for many of the tasks we want LLMs to perform. This warning also applies to context-specific modifications to the model such as alignment and fine-tuning: strategies that do not attempt to reconcile the innate gnogeographic constraints of the model with the intended task have a flimsy foundation on which to build success. As long as generating plausible text is the goal for the model, we should expect plausible text rather than accurate text to be the output. That those two can often coincide, while alluring and often convenient, is fundamentally incidental.122122122We think this framework for thinking about LLMs can be helpful to people trying to decide how to use and study them. If an LLM is deployed as part of e.g., a transcription pipeline, the potential for unwanted hallucination can and should be assumed by default Koenecke et al. (2024), unless part of the pipeline’s design is specifically intended to counteract not just the symptomatic output but the causes - and even then, caution is necessary.
@@ 5 Post-Script: what’s next
The architectural choice of the token, while foundational, may be on its way out for some AI applications. By hardwiring a specific, language-centric unit of cognition, tokens impose a form of instantiated linguistic relativism – a commitment to the primacy of language for thought. As discussed above, the tokens themselves are determinative of distributional and organizational unit, and carry consequences such as ‘flattening’. However, the commitment to linguistic primacy is even more pervasive – it manifests not only in tokens’ role as semantic primitives and as the units governing the distributional information perceptible to the model, but also in their function in Chain-of-Thought-like (CoT) reasoning123123123What exactly ‘reasoning’ models are doing – whether via CoT or other means – remains an open question. We believe there are compelling reasons to think the gnogeographic ceiling of deep learning could be far higher than that attested by current prominent LLMs Zimmerman (2025). However, model behavior is notoriously difficult to interpret, so caution is warranted. For concision, here we assume CoT induces something beyond MVP linguistic performance – a substantive rather than stylistic difference. However, it is possible that ‘reasoning’ LLMs may more aptly be described as reweighting statistical patterns under varying conditions rather than truly reasoning about new tasks. and their direct relationship with computational resource allocation during inference (e.g., the correspondence between a single forward pass and a single token during generation) Gladstone et al. (2025); Wang et al. (2025). Historically, more open-ended architectures have often unlocked greater performance, suggesting that the degree of cognition currently outsourced to tokens as an architectural choice might be better left to the model itself to learn Zimmerman (2025).
@@
We speculate that the answer lies in the expanding scope of knowledge required by LLM tasks and thus incorporated into LLM training objectives. NTP more-or-less suffices for human-like text generation. However, the shift towards ‘reasoning’ – epitomized by CoT prompting – demands a broader scope. The task extends beyond MVP linguistic competence to solving multipart logical problems that invoke narrative, social rules, specialized techniques, and extensive world knowledge. Consequently, effective CoT and instruction tuning must engage larger linguistic structures than NTP alone. We view CoT as a possible extension of the NTP objective, broadening the model’s perceptible range of knowledge and raising the ceiling of its gnogeography.124124124Essentially, as tasks demand more from the information encoded in language, models must become increasingly aware of linguistic structures not salient within NTP alone (at least generally, many of these structures would seem to be larger than that local scope). Thus, as ‘reasoning’ LLMs improve, they increasingly demonstrate the feasibility of language as a cognitive foundation, one strategy for how it is possible to think. Caveats remain: the boundary between world knowledge and linguistic knowledge is unclear, perhaps non-existent, and NTP may foster some non-local abstractions Dumas et al. (2025); Cui et al. (2024); Zimmerman (2025); Hewitt et al. (2024); Digutsch and Kosinski (2023); Huang et al. (2023b). Relationships among task design, implementation, model capacity, and data remain only loosely understood – but in broad strokes, this is the emerging picture.
@@
Simultaneously, in refining what they want LLMs to do, computer scientists are often rediscovering principles from linguistic pragmatics. For instance, the CoT principles of efficacy, efficiency, brevity, and sufficiency directly echo Grice’s maxims of quantity, quality, relation, and manner Cheng et al. (2025); Grice (1989). Similarly, a Gricean logic is evident in efforts to incentivize ‘truthfulness’ by teaching a model to abstain when uncertain Wei et al. (2025),125125125‘Truthfulness’ is a loaded concept mostly out of scope for this discussion; we note ‘truth’ is – at the very least – unlikely to be a property found within the four corners of the text, loosely paraphrasing Joseph Weizenbaum. and in instruction tuning. These maxims emerge at the scale of conversation and multi-utterance discourse – scales that ‘reasoning’ models are increasingly forced to navigate. These resonances demonstrate that computer scientists could look to linguistics (and psychology and cognitive science) for inspiration, and in particular, for strategies which could enhance model performance – and vice versa!126126126Regardless of their architectural suitability, tokens remain invaluable to linguists: they are configurable, language-related structures generated by the most performant computational models of language yet developed. We plan to explore their use in linguistics further in future work (and hope others will be doing the same)!
@@
A growing body of work lends support to our arguments herein (first shared in preprint, December 2024). Several recent works support this view of tokens as foundational semantic and distributional primitives Oh and Schuler (2024). Gurnee and Tegmark (2024) show that learned spatiotemporal representations are robust to most prompting changes, yet performance degrades with random distracting tokens or atypical capitalizations, supporting the view that mapping a concept across multiple tokenizations obscures its unity for the model (see also Figure 10). Marjieh et al. (2025) show that conceptual entanglement arises from dual string-digit representations of numbers, leading to a “string bias” that can result in incorrect answers. Along these same lines, see discussion of ‘evil numbers’ in Betley et al. (2025). With respect to the role of the NTP objective function, McCoy et al. (2024a) likewise argue that deep learning models should be understood through their objective functions,127127127They term this the ‘teleological approach’; we referred to it as ‘POSIWID teleology’ after Stafford Beer., that such models often struggle with atypical data, and that the conflation of frequency and truth, though powerful, carries substantial risks.
@@
Finally, Cui et al. (2024) identify a phase transition we interpret as marking a shift between pattern-matching (regime 1) and abstraction-based (regime 2) learning.128128128Consistent with our Discussion (Sec. 3), and further developed in Zimmerman (2025)., 129129129In conjunction, recall the earlier quote from Digutsch and Kosinski (2023): “GPT-3’s semantic activation is better predicted by similarity in words’ meaning (i.e., semantic similarity) rather than their co-occurrence in the language (i.e., associative similarity)”. Their results suggest that while LLMs embody a commitment to token-based linguistic primacy in cognition – with its attendant token-structure-based pitfalls – models may be able to build abstractions atop that substrate, so that, practically, they may not be overly constrained by that choice. This success is monumental: that LLMs can produce output that looks like ‘thinking’ through a cognition with (a form of) language baked into its core is revelatory for multiple fields (including philosophy, AI, and linguistics).
@@
However, it also has profound implications for LLM ‘experience’. LLMs, with such curtailed bodies and access to the world as we know it, experience the world as and through text, which appears to them nihilogonously. As a result, all words are speech acts, and all actions are speech – an extreme, perhaps degenerative,130130130Maybe even perverse. instantiation of the idea that “all thoughts are thinkers” Levin (2024). Ultimately, this defines the gnogeographic ceiling of pure LLMs, and may continue to echo through their descendant iterations.131131131In other words, LLMs have – excitingly! – built abstract thought on a language-only foundation, but this success also traps them in a reality made entirely out of words, which fundamentally limits what they can be. This idea is further developed in Zimmerman et al. (2024)/ Zimmerman (2025).

### v9

@@ 4 Conclusion
It is clear that we need to develop strategies that allow LLMs to make better use of the data to which they are exposed.113113113Zhuang et al., 2024 note “the immense volume of training data necessitated for effective LM performance, surpassing—by orders of magnitude—the linguistic input received during human language acquisition” Zhuang, Fedorenko, and Andreas (2024). People, by contrast, are very good at wringing new information from familiar stories.114114114Right now, models display a striking dissimilarity with people in what happens when they are repeatedly exposed to the same data, either directly, or indirectly through iterated training on synthetic data. For example, Chang et al., 2024 found that “LLMs trained with duplicated training data exhibit faster forgetting” Chang et al. (2024), and Shumailov et al., 2024, found “indiscriminate use of model-generated content in training causes irreversible defects in the resulting models, in which tails of the original content distribution disappear” Shumailov et al. (2024). On the other hand, a hallmark of human storytelling is enjoying—and getting new content out of—experiencing a story (created by other humans, even potentially by oneself) again and again. People have been studying religious texts, as in Biblical exegesis, for thousands of years, yet new ideas emerge. An individual may return to the Bible (continuing the same example) upwards of hundreds of times, and yet experience new insight. How can LLMs wring new insight from old data? This stark difference in behaviour is worth further consideration.
@@
One strategy would be to bring various representations the LLM uses into greater alignment with human structures and processes. There is evidence that humans can do quite a lot with very little data (e.g., a broad spectrum of poverty-of-the-stimulus-related arguments and processes Zhuang, Fedorenko, and Andreas (2024), including fast-mapping Gelman and Brandone (2010), and in experimental scenarios testing the distributional hypothesis McDonald and Ramscar (2001)).115115115Likewise, LLMs can sometimes do a lot with a little, such as in lexinvariant LMs Huang et al. (2023b).
@@
Furthermore, good model vocabularies may already be converging on being somewhat human-like. Symmetrically, the tokens themselves point to commonalities between the entropy-minimizing objective functions often used in tokenization and the forces that shape human language. The hundreds of thousands of years Fedorenko, Piantadosi, and Gibson (2024)116116116E. Fedorenko is an author on both Fedorenko et al.,, 2024 Fedorenko, Piantadosi, and Gibson (2024) and Zhuang et al.,, 2024 Zhuang, Fedorenko, and Andreas (2024), which seem to make arguments that are at least partially contradictory. We assume in good faith there must be a perspective (presumably E.F.’s) from which the arguments are entirely consistent, and take both papers at face value for the purposes of this project. in which only people spoke language, during which pressures of efficiency seem to have shaped its development at multiple levels, surely evolved structures we can now take advantage of in our models.
@@
We argue that the de facto approaches in each of those areas are worth revisiting.117117117The use of special tokens (e.g., SEP) in LLMs is out of scope for this paper but intriguing for future work, as is the impact of embedding vector dimension and other implementation choices. More broadly, now that the efficacy of LLM architecture has been established, we should revisit our prior underconstrained implementation choices, such as the relevant objective functions that shape the model and the functions that conjoin its components.118118118e.g., operations involving Query, Key, and Value vectors in the attention blocks, or applications of softmax.
@@
We contend that LLMs validate and isolate the distributional hypothesis. Specifically, they show the power of syntagmatic information for producing minimally viable language performance (when coupled with sufficient architecture). We hypothesize that LLMs bootstrap limited meaning, potentially including abstractions and world knowledge, from distributional patterns: more specifically, that while LLM’s start with syntagmatics, they create further structure leading to limited semantic meaning, the ceiling of which is what can be conveyed by purely distributional linguistic information and diegetic approximations composed thereof.119119119Harris’ purely linguistic meaning Sahlgren (2008). The lack of hard boundary between world knowledge and linguistic knowledge makes this feat possible but it also engenders the problems that come from fuzzing up grammaticality and truth, or frequency and truth Zimmerman et al. (2024); Gong et al. (2020),120120120Gong et al., 2020, found that in some situations, word embeddings were “biased towards word frequency”, with the embeddings of high- and low-frequency words lying in different areas of the embedding space, and the embedding of a rare word and a popular word being physically distant despite their semantic similarity. Within the embedding space, rare and popular words behaved differently, with popular words having semantically relevant neighbors, and rare words being overwhelmingly surrounded by other, semantically-unrelated, rare words Gong et al. (2020). (They foresee these structures as causing downstream performance issues, so they propose FRequency AGnostic word Embedding (FRAGE) as a solution Gong et al. (2020).) which are amplified by the choice of next token prediction as objective function.121121121And by other, more specific, architectural choices, such as how attention is implemented. These choices are related, since the attention design is in the service of generating plausible output tokens, so in the context of this paper, we are referring to the broad goal of getting the model to generate plausible tokens, not only to specific portions of the model that might be most directly technically implicated, like the decoder (if there is one), the output layer, or the final probability distribution over the vocabulary. We propose that experiential grounding enhances the depth of meaning through othermodal reference for both people and LLMs122122122Recall Wolfe and Caliskan, 2022 Wolfe and Caliskan (2022) and Zhuang et al., 2024 Zhuang, Fedorenko, and Andreas (2024). and therefore represents a promising avenue for their continued improvement.
@@
Importantly, a model’s gnogeography123123123Recall that gnogeography is the landscape of knowledge; a being’s mind-and-body, their architecture, in conjunction with what information they have taken in. can be expanded through diegetic approximations Zimmerman et al. (2024); Brown et al. (2020); Chang et al. (2024); Mahowald et al. (2024). Meaning that, with careful structuring of training data and model architecture (e.g., batch size, number of training steps), LLMs do seem to be able to acquire additional world knowledge Chang et al. (2024), implying that there is space between the world knowledge necessary for MVP language and what it is possible to know through the medium of language. Given that current LLMs clearly possess MVP language but are missing much human-familiar knowledge and common sense, they are somewhere within that space Zimmerman et al. (2024); Mahowald et al. (2024). The entirety of that space is the upper bound of their gnogeography.
This is crucial to remember in the face of AI hype, one core promise of which is “we have a machine that can talk, therefore we have a machine that can think”. In fact, LLMs demonstrate the plausibility of a machine that can solely talk; while a machine may be able to both talk and think, the ‘therefore’ between those abilities is not justified.124124124“Talk” and “think” are shorthand. Of course, talking involves thinking, and thinking can be scaffolded by talking. See previous sections, especially Sec. 2.3, where we referenced stochastic parrots; we do think MVP language as implemented in LLMs involves more understanding than would be granted by that framework.
The size and shape of the upper bound LLM gnogeography is unknown.125125125The hypothetical ceiling being everything that can be encoded in language. While this is exciting, it also means it is plausible that their gnogeography is incompatible with what would be required for many of the tasks we want LLMs to perform. This warning also applies to context-specific modifications to the model such as alignment and fine-tuning: strategies that do not attempt to reconcile the innate gnogeographic constraints of the model with the intended task have a flimsy foundation on which to build success. As long as generating plausible text is the goal for the model, we should expect plausible text rather than accurate text to be the output. That those two can often coincide, while alluring and often convenient, is fundamentally incidental.126126126We think this framework for thinking about LLMs can be helpful to people trying to decide how to use and study them. If an LLM is deployed as part of e.g., a transcription pipeline, the potential for unwanted hallucination can and should be assumed by default Koenecke et al. (2024), unless part of the pipeline’s design is specifically intended to counteract not just the symptomatic output but the causes - and even then, caution is necessary.
@@ 5 Post-Script: what’s next
The architectural choice of the token, while foundational, may be on its way out for some AI applications. By hardwiring a specific, language-centric unit of cognition, tokens impose a form of instantiated linguistic relativism – a commitment to the primacy of language for thought. As discussed above, the tokens themselves are determinative of distributional and organizational unit, and carry consequences such as ‘flattening’. However, the commitment to linguistic primacy is even more pervasive – it manifests not only in tokens’ role as semantic primitives and as the units governing the distributional information perceptible to the model, but also in their function in Chain-of-Thought-like (CoT) reasoning127127127What exactly ‘reasoning’ models are doing – whether via CoT or other means – remains an open question. We believe there are compelling reasons to think the gnogeographic ceiling of deep learning could be far higher than that attested by current prominent LLMs Zimmerman (2025). However, model behavior is notoriously difficult to interpret, so caution is warranted. For concision, here we assume CoT induces something beyond MVP linguistic performance – a substantive rather than stylistic difference. However, it is possible that ‘reasoning’ LLMs may more aptly be described as reweighting statistical patterns under varying conditions rather than truly reasoning about new tasks. and their direct relationship with computational resource allocation during inference (e.g., the correspondence between a single forward pass and a single token during generation) Gladstone et al. (2025); Wang et al. (2025). Historically, more open-ended architectures have often unlocked greater performance, suggesting that the degree of cognition currently outsourced to tokens as an architectural choice might be better left to the model itself to learn Zimmerman (2025).
@@
We speculate that the answer lies in the expanding scope of knowledge required by LLM tasks and thus incorporated into LLM training objectives. NTP more-or-less suffices for human-like text generation. However, the shift towards ‘reasoning’ – epitomized by CoT prompting – demands a broader scope. The task extends beyond MVP linguistic competence to solving multipart logical problems that invoke narrative, social rules, specialized techniques, and extensive world knowledge. Consequently, effective CoT and instruction tuning must engage larger linguistic structures than NTP alone. We view CoT as a possible extension of the NTP objective, broadening the model’s perceptible range of knowledge and raising the ceiling of its gnogeography.128128128Essentially, as tasks demand more from the information encoded in language, models must become increasingly aware of linguistic structures not salient within NTP alone (at least generally, many of these structures would seem to be larger than that local scope). Thus, as ‘reasoning’ LLMs improve, they increasingly demonstrate the feasibility of language as a cognitive foundation, one strategy for how it is possible to think. Caveats remain: the boundary between world knowledge and linguistic knowledge is unclear, perhaps non-existent, and NTP may foster some non-local abstractions Dumas et al. (2025); Cui et al. (2024); Zimmerman (2025); Hewitt et al. (2024); Digutsch and Kosinski (2023); Huang et al. (2023b). Relationships among task design, implementation, model capacity, and data remain only loosely understood – but in broad strokes, this is the emerging picture.
@@
Simultaneously, in refining what they want LLMs to do, computer scientists are often rediscovering principles from linguistic pragmatics. For instance, the CoT principles of efficacy, efficiency, brevity, and sufficiency directly echo Grice’s maxims of quantity, quality, relation, and manner Cheng et al. (2025); Grice (1989). Similarly, a Gricean logic is evident in efforts to incentivize ‘truthfulness’ by teaching a model to abstain when uncertain Wei et al. (2025),129129129‘Truthfulness’ is a loaded concept mostly out of scope for this discussion; we note ‘truth’ is – at the very least – unlikely to be a property found within the four corners of the text, loosely paraphrasing Joseph Weizenbaum. and in instruction tuning. These maxims emerge at the scale of conversation and multi-utterance discourse – scales that ‘reasoning’ models are increasingly forced to navigate. These resonances demonstrate that computer scientists could look to linguistics (and psychology and cognitive science) for inspiration, and in particular, for strategies which could enhance model performance – and vice versa!130130130Regardless of their architectural suitability, tokens remain invaluable to linguists: they are configurable, language-related structures generated by the most performant computational models of language yet developed. We plan to explore their use in linguistics further in future work (and hope others will be doing the same)!
@@
A growing body of work lends support to our arguments herein (first shared in preprint, December 2024). Several recent works support this view of tokens as foundational semantic and distributional primitives Oh and Schuler (2024). Gurnee and Tegmark (2024) show that learned spatiotemporal representations are robust to most prompting changes, yet performance degrades with random distracting tokens or atypical capitalizations, supporting the view that mapping a concept across multiple tokenizations obscures its unity for the model (see also Figure 10). Marjieh et al. (2025) show that conceptual entanglement arises from dual string-digit representations of numbers, leading to a “string bias” that can result in incorrect answers. Along these same lines, see discussion of ‘evil numbers’ in Betley et al. (2025). With respect to the role of the NTP objective function, McCoy et al. (2024a) likewise argue that deep learning models should be understood through their objective functions,131131131They term this the ‘teleological approach’; we referred to it as ‘POSIWID teleology’ after Stafford Beer., that such models often struggle with atypical data, and that the conflation of frequency and truth, though powerful, carries substantial risks.
@@
Finally, Cui et al. (2024) identify a phase transition we interpret as marking a shift between pattern-matching (regime 1) and abstraction-based (regime 2) learning.132132132Consistent with our Discussion (Sec. 3), and further developed in Zimmerman (2025)., 133133133In conjunction, recall the earlier quote from Digutsch and Kosinski (2023): “GPT-3’s semantic activation is better predicted by similarity in words’ meaning (i.e., semantic similarity) rather than their co-occurrence in the language (i.e., associative similarity)”. Their results suggest that while LLMs embody a commitment to token-based linguistic primacy in cognition – with its attendant token-structure-based pitfalls – models may be able to build abstractions atop that substrate, so that, practically, they may not be overly constrained by that choice. This success is monumental: that LLMs can produce output that looks like ‘thinking’ through a cognition with (a form of) language baked into its core is revelatory for multiple fields (including philosophy, AI, and linguistics).
@@
However, it also has profound implications for LLM ‘experience’. LLMs, with such curtailed bodies and access to the world as we know it, experience the world as and through text, which appears to them nihilogonously. As a result, all words are speech acts, and all actions are speech – an extreme, perhaps degenerative,134134134Maybe even perverse. instantiation of the idea that “all thoughts are thinkers” Levin (2024). Ultimately, this defines the gnogeographic ceiling of pure LLMs, and may continue to echo through their descendant iterations.135135135In other words, LLMs have – excitingly! – built abstract thought on a language-only foundation, but this success also traps them in a reality made entirely out of words, which fundamentally limits what they can be. This idea is further developed in Zimmerman et al. (2024)/ Zimmerman (2025).